OpenStudy (anonymous):

Solve for x: |2x + 10| = 18 x = −4, x = 14 x = 4, x = −4 x = 14, x = −14 x = 4, x = −14

OpenStudy (anonymous):

Its B

OpenStudy (zzr0ck3r):

no its not @GodsGirl19

OpenStudy (zzr0ck3r):

|2x+10|=18 so 2x+10=18 and 2x+10=-18 x=4,x=-14

OpenStudy (anonymous):

my bad I forgot to do the 2x+10=-18 sorry

Directrix (directrix):

@GodsGirl19 Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

why a lot of other people do it